# Photonuclear Production of 195mPt for Medical Applications: Cross Sections and Energy Thresholds

### v1

@@ Abstract
Platinum radioisotopes are of growing interest for targeted cancer therapy and diagnostic imaging because their decay delivers highly localized radiation doses in tissue, herewith enabling precise DNA damage through Auger-electron emission. Developing production technologies that provide platinum isotopes with high specific activity is therefore essential, and photonuclear reactions on stable nuclei offer a viable accelerator-based route when supported by reliable cross-section data. We report photonuclear cross-section measurements for the 197Au(,pn)195mPt reaction at incident -ray energies of 27, 29, and 31 MeV using the activation method. The measurements were performed by irradiating a stack of concentric-ring gold targets with a quasi-monoenergetic -ray beam provided by the High Intensity Gamma-ray Source (HIS). The induced 195mPt activity was quantified using off-line -ray spectroscopy. The present results provide the first experimental information in the energy region where the 197Au(,pn)195mPt reaction becomes significant and demonstrate the minimum energy requirements and feasibility of producing 195mPt via photonuclear reactions using electron accelerator facilities. These measurements indicate that electron energies of 50 MeV are required to achieve practically meaningful production yields.
@@ I Introduction
Platinum-195m (195mPt) is a metastable radionuclide with a half-life of about 4 days that decays via isomeric transitions. Because of its highly internally converted decay, 195mPt emits a large number of low-energy Auger electrons, which deposit their energy over nanometer-scale distances and have attracted interest for targeted Auger-electron therapy and theranostic applications [6, 9]. In addition, the emission of a 98.9-keV -ray makes 195mPt suitable for quantitative SPECT imaging [1]. The medical relevance of 195mPt is closely linked to platinum-based chemotherapy. Platinum compounds, such as cisplatin, are widely used anticancer drugs, and radiolabeling these agents with 195mPt enables direct imaging of their biodistribution and pharmacokinetics without altering their chemical structure. Preclinical studies have demonstrated the feasibility of SPECT imaging and quantitative analysis using 195mPt-labeled compounds [1, 3], and more recent translational work has further highlighted the potential of 195mPt for imaging chemotherapy delivery in vivo [5].
Production of 195mPt has traditionally relied on reactor-based routes involving neutron capture on enriched iridium targets [10]. Although these methods can provide substantial yields, they may be limited in achievable specific activity due to the presence of stable platinum isotopes. Accelerator-based production using charged particles, such as the 192Os(,n)195mPt reaction, has been shown to improve specific activity, but typically requires enriched targets and is constrained by available beam currents [4]. Production of 195mPt via photonuclear reactions using electron accelerators has received comparatively little experimental attention [2]. This approach is attractive because photonuclear reactions can produce the desired radionuclide in a chemically distinct target material, herewith enabling efficient chemical separation and potentially high specific activity, an important requirement for Auger-electron-emitting radionuclides [9]. For feasibility studies of photonuclear production of 195mPt, reliable cross-section data for the relevant reactions are required. For 195mPt-based radiopharmaceuticals, the effective specific activity is governed, not only by the total platinum content, but also by the isomer-to-ground-state production ratio. Theoretical predictions based on TALYS [8] indicate that this ratio remains 1 for photon energies above 30 MeV.
@@ III Experiment
A stack of five targets, each divided into three segments [12], was irradiated at seven different central photon beam energies at the High Intensity Gamma-ray Source (HIS) . Owing to the radial energy gradient across the target stack, the photon energy at the outermost ring was typically close to the central energy of the subsequent lower-energy irradiation, providing an internal consistency check for the measurements (see ref. [12] for details). Using targets composed of three concentric segments, cross sections for the 197Au(,pn)195mPt reaction were obtained at a total of 3 distinct photon energies for the 31 MeV run, as the cross sections are extremely low below 30 MeV. Each irradiation was performed for 8–12 hours using a photon beam with fluxes in the range of - . Detailed information on incident beam flux, irradiation time, target thickness, and beam energy and flux at each segment is summarized in Tables III, IV, and VIII of Ref. [12]. The measurement and cooling times for 31 MeV run are summarized in Table 1.
@@
The expanded spectra around the 98.9-keV region (see Figs. 2 (b)–(d)) demonstrate a clear evolution of the relative peak intensities with cooling time. At early times after irradiation, the spectrum is dominated by contributions from shorter-lived nuclides, whereas measurements performed after longer cooling times increasingly emphasize the decay of longer-lived reaction products. This temporal variation arises from the different half-lives of 195mPt ( d) and 195Au ( d), and enables their individual activities to be disentangled through measurements performed at multiple cooling times. The systematic change in the 98.9-keV peak intensity as a function of cooling time provides the experimental basis of the approach for the activity decomposition described in Sec. II. By combining the spectra measured at different cooling times with a least-squares analysis, the initial activities of the contributing nuclides can be extracted reliably, despite the spectral overlap at this 98.9-keV -ray energy.
@@ V Conclusion
We have measured the photonuclear cross sections for the 197Au(,pn)195mPt reaction using a quasi-monoenergetic -ray beam at incident energies of 27, 29, and 31 MeV at HIS. The observed 98.9-keV -ray peak contains contributions from the decays of both 195Au and 195mPt. These contributions were successfully extracted using a least-squares analysis of measurements performed at multiple cooling times. For the first time, cross sections were experimentally determined in the vicinity of the reaction threshold. No statistically significant 98.9-keV peak attributable to 195mPt was observed for irradiation energies below 27 MeV, despite the reaction threshold energy of Eth = 13.96 MeV, indicating that, under present experimental conditions, specifically a photon flux of and irradiation time 8-10 hours, the production yield is insufficient to detect this rare reaction channel at lower energies. The measured cross sections indicate that the 197Au(,pn)195mPt reaction begins to contribute measurably at incident photon energies of 30 MeV, but remains too weak for practical isotope production with bremsstrahlung rays at energies typically available at electron accelerator facilities. Although electron energies around 40 MeV are commonly employed for photonuclear isotope production, the present results suggest that such energies are insufficient to yield meaningful 195mPt production under realistic experimental conditions. Combining our results with the model predictions for the excitation function for the 197Au(,pn)195mPt reaction shown in Fig. 3, we project that efficient production of 195mPt via photonuclear reactions in quantities required for applications will require bremsstrahlung photon beams with substantially higher end-point energies than are standard, e.g., on the order of 50-60 MeV. The present results provide experimental constraints on the minimum photon energies relevant for photonuclear production of 195mPt and motivate future measurements at higher -ray energies, extending beyond 50 MeV at HIS.

### v2

# Photonuclear Cross Sections for the 197Au(,pn)195mPt Reaction Near Threshold

@@ Abstract
Platinum radioisotopes are of growing interest for targeted cancer therapy and diagnostic imaging because their decay delivers highly localized radiation doses in tissue, herewith enabling precise DNA damage through Auger-electron emission. Developing production technologies that provide platinum isotopes with high specific activity is therefore essential, and photonuclear reactions on stable nuclei offer a viable accelerator-based route when supported by reliable cross-section data. We report photonuclear cross-section measurements for the 197Au(,pn)195mPt reaction at incident -ray energies of 27, 29, and 31 MeV using the activation method. The measurements were performed by irradiating a stack of concentric-ring gold targets with a quasi-monoenergetic -ray beam provided by the High Intensity Gamma-ray Source (HIS). The induced 195mPt activity was quantified using off-line -ray spectroscopy. These data provide the first experimental constraints on the 197Au(,pn)195mPt cross section in the near-threshold region. The measured excitation function is compared with PHITS and TALYS calculations and indicates that the reaction becomes measurable only near 30 MeV, with substantially higher bremsstrahlung end-point energies required for practically meaningful production.
@@ I Introduction
Platinum-195m (195mPt) is a metastable radionuclide with a half-life of about 4 days that decays via isomeric transitions. Because of its highly internally converted decay, 195mPt emits a large number of low-energy Auger electrons, which deposit their energy over nanometer-scale distances and have attracted interest for targeted Auger-electron therapy and theranostic applicationsΒ [6, 9]. In addition, the emission of a 98.9-keV -ray makes 195mPt suitable for quantitative SPECT imagingΒ [1]. The medical relevance of 195mPt is closely linked to platinum-based chemotherapy. Platinum compounds, such as cisplatin, are widely used anticancer drugs, and radiolabeling these agents with 195mPt enables direct imaging of their biodistribution and pharmacokinetics without altering their chemical structure. Preclinical studies have demonstrated the feasibility of SPECT imaging and quantitative analysis using 195mPt-labeled compoundsΒ [1, 3], and more recent translational work has further highlighted the potential of 195mPt for imaging chemotherapy delivery in vivoΒ [5].
Production of 195mPt has traditionally relied on reactor-based routes involving neutron capture on enriched iridium targetsΒ [10]. Although these methods can provide substantial yields, they may be limited in achievable specific activity due to the presence of stable platinum isotopes. Accelerator-based production using charged particles, such as the 192Os(,n)195mPt reaction, has been shown to improve specific activity, but typically requires enriched targets and is constrained by available beam currentsΒ [4]. Production of 195mPt via photonuclear reactions using electron accelerators has received comparatively little experimental attention [2]. This approach is attractive because photonuclear reactions can produce the desired radionuclide in a chemically distinct target material, herewith enabling efficient chemical separation and potentially high specific activity, an important requirement for Auger-electron-emitting radionuclides [9]. For feasibility studies of photonuclear production of 195mPt, reliable cross-section data for the relevant reactions are required. For 195mPt-based radiopharmaceuticals, the effective specific activity is governed, not only by the total platinum content, but also by the isomer-to-ground-state production ratio. Theoretical predictions based on TALYS [8] indicate that this ratio remains 1 for photon energies above 30 MeV.
@@ III Experiment
A stack of five targets, each divided into three segments [12], was irradiated at seven different central photon beam energies at the High Intensity Gamma-ray Source (HIS) . Owing to the radial energy gradient across the target stack, the photon energy at the outermost ring was typically close to the central energy of the subsequent lower-energy irradiation, providing an internal consistency check for the measurements (see ref. [12] for details). Using targets composed of three concentric segments, cross sections for the 197Au(,pn)195mPt reaction were obtained at a total of 3 distinct photon energies for the 31 MeV run, as the cross sections are extremely low below 30 MeV. Each irradiation was performed for 8β12 hours using a photon beam with fluxes in the range of - . Detailed information on incident beam flux, irradiation time, target thickness, and beam energy and flux at each segment is summarized in Tables III, IV, and VIII of Ref. [12]. The measurement and cooling times for 31 MeV run are summarized in Table 1.
@@
The expanded spectra around the 98.9-keV region (see Figs. 2 (b)β(d)) demonstrate a clear evolution of the relative peak intensities with cooling time. At early times after irradiation, the spectrum is dominated by contributions from shorter-lived nuclides, whereas measurements performed after longer cooling times increasingly emphasize the decay of longer-lived reaction products. This temporal variation arises from the different half-lives of 195mPt ( d) and 195Au ( d), and enables their individual activities to be disentangled through measurements performed at multiple cooling times. The systematic change in the 98.9-keV peak intensity as a function of cooling time provides the experimental basis of the approach for the activity decomposition described in Sec. II. By combining the spectra measured at different cooling times with a least-squares analysis, the initial activities of the contributing nuclides can be extracted reliably, despite the spectral overlap at this 98.9-keV -ray energy.
@@ V Conclusion
We have measured the photonuclear cross sections for the 197Au(,pn)195mPt reaction at incident energies of 27, 29, and 31 MeV using a quasi-monoenergetic -ray beam at HIS. The observed 98.9-keV -ray peak contains contributions from the decays of both 195Au and 195mPt. These contributions were successfully extracted using a least-squares analysis of measurements performed at multiple cooling times. For the first time, cross sections were experimentally determined in the vicinity of the reaction threshold. No statistically significant 98.9-keV peak attributable to 195mPt was observed for irradiation energies below 27 MeV, despite the reaction threshold energy of Eth = 13.96 MeV, indicating that, under present experimental conditions, specifically a photon flux of and irradiation time 8-10 hours, the production yield is insufficient to detect this rare reaction channel at lower energies. The measured cross sections indicate that the 197Au(,pn)195mPt reaction begins to contribute measurably at incident photon energies of 30 MeV, and provide the first experimental constraints on the cross section in the near-threshold region. The measured cross sections are consistent, within the experimental uncertainties, with the overall trends predicted by the PHITS and TALYS calculations. Although electron energies 30β40 MeV are commonly employed for photonuclear isotope production, the present results suggest that such energies are insufficient to yield meaningful 195mPt production under realistic experimental conditions. Combining our results with the model predictions for the excitation function for the 197Au(,pn)195mPt reaction shown in Fig. 3, we project that efficient production of 195mPt via photonuclear reactions in quantities required for applications will require bremsstrahlung photon beams with substantially higher end-point energies than are standard, e.g., on the order of 50-60 MeV. The present results provide new benchmark data for calculations of photonuclear production of 195mPt and motivate future measurements at higher -ray energies, extending beyond 50 MeV at HIS.